OpenStudy (anonymous):

Simplify 2 cubed root of 128

OpenStudy (anonymous):

128 = 2*64 = 2*8^2 \[2\sqrt{128} = 2\sqrt{2*8^{2}}= 16\sqrt{2}\]

OpenStudy (anonymous):

that not one of the answers

OpenStudy (anonymous):

\[2\left(128^{1/3}\right)=8 \left(2^{1/3}\right) \]

OpenStudy (anonymous):

opps... \[2\sqrt[3]{128} = 2\sqrt[3]{2*8^{2}}= 2\sqrt[3]{2*2^{3^{2}}} = 8\sqrt[3]{2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.